OpenStudy (jmartinez638):

A 675 kg train car coasts slowly (v = 2.4m/sec) under a hopper that fills it with coal. If the car's velocity slows to 1 m/sec, what is the combined mass of the train car and the coal?

OpenStudy (owen3):

The rate of coal entering from the hopper is not given?

OpenStudy (jmartinez638):

No. Is the question broken?

OpenStudy (owen3):

maybe not :)

OpenStudy (jmartinez638):

I think the hopper is irrelevant. So the question would be:

OpenStudy (owen3):

I think we have to assume the addition of coal slows down the train.

OpenStudy (jmartinez638):

The train car weighs 675 kg and moves at a velocity of 2.4m/sec while empty, and with the addition of coal slows down to 1 m/sec.

OpenStudy (jmartinez638):

Is that right?

OpenStudy (owen3):

correct, so im thinking we can use momentum here somehow

OpenStudy (owen3):

can we assume conservation of momentum. in which case we have p1 = p2 m1*v1 = m2*v2 675 * 2.4 = m2 * 1

OpenStudy (owen3):

which evaluates to 1620 kg. that seems too easy though

OpenStudy (jmartinez638):

Hmmm. It does, doesn't it?

OpenStudy (owen3):

we should add some more details before the coal enters the train, we have a system of two masses. the train and the coal. the train moves at velocity v1 = 2.4 the coal has velocity zero (before it enters the train). After the coal gets loaded we have a combined mass m1 + m2 which moves at velocity v final p1 before = p2 after m1v1 + m2 v2 = (m1 + m2) * v 675 * 2.4 + m2 * 0 = (675 + m2) * 1 from which we derive m1 + m2 = 1620 kg so that is a better solution

OpenStudy (jmartinez638):

That makes more sense. Sorry for the delay, I got distracted for a bit.